OpenStudy (marcelie):

help please !!!! Find the dimensions of the rectangular corral producing the greatest enclosed area given 200 feet of fencing.

imqwerty (imqwerty):

ok so its like- ur given that- x+y=200 y=200-x and the area should be maximum area=x*y area=x*(200-x) area=200x-x^2 u can get the value of x for which the area will be max by differentiating the area equation which we got with respect to x and equating it to 0 so u get \[\frac{ d}{ dx }(200x-x^2) =0\] after differentiating u get- 200-2x=0 200=2x x=100 and we know that y=200-x so y=200-100 y=100 so the dimensions are x=100 and y=100 so basically if u are maximizing the area then the figure has to become a square even tho we are given that its a rectangle :)